# algorithmic

### v1

@@ 1 Introduction
Recently, deep learning-based approaches have gained attention for overcoming the limitations of traditional graph generation methods by learning the complex topology of real-world graphs. These approaches typically involve an encoder that learns a dense representation of the graph, a sampler that generates samples from this representation, and a decoder which restores the sampled representation into a graph structure (Zhu et al., 2022). Different decoding strategies have been proposed, including sequential decoders that generate the graph step-by-step (You et al., 2018; Liao et al., 2019), and one-shot generators that can create the entire graph in a single step (Kipf & Welling, 2016b; Simonovsky & Komodakis, 2018; Grover et al., 2019). While sequential decoders struggle to model long-term dependencies and require node-ordering process, one-shot generators suffer from scalability issues and compromised quality due the to treatment of edges as independent (Chanpuriya et al., 2021), making these methods suboptimal for large-scale graph generation (Guo & Zhao, 2022).
@@
An even more recent body of work in graph generation are diffusion-based probabilistic models, inspired by non-equilibrium thermodynamics and first introduced by Sohl-Dickstein et al. (2015). Briefly, diffusion models consist of two processes: A forward diffusion process which gradually corrupts input data until it reaches pure noise, and a reverse diffusion process which learns the generative mechanism of the original input data using a neural network. Diffusion-based methods for graph generation can be divided into two main categories. The first one includes methods that implement diffusion in the continuous space e.g., by adding Gaussian noise to the node features and graph adjacency matrix (Niu et al., 2020; Jo et al., 2022). This form of diffusion however makes it difficult to capture the underlying structure of graphs since it destroys their sparsity pattern (Vignac et al., 2023). The second category includes methods that are based on diffusion in the discrete space (Haefeli et al., 2022; Vignac et al., 2023; Chen et al., 2023) by performing successive graph modifications e.g., adding or deleting edges/nodes or edge/node features. Diffusion-based graph generation methods are invariant to node ordering and do not suffer from long-term memory dependency which makes them advantageous over autoregressive (sequential) methods. However, many approaches found in the literature are only designed for small graphs (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2023). One of the most recent diffusion-based graph generation approaches is EDGE (Chen et al., 2023), which uses a combination of a discrete diffusion model and a Graph Neural Network (GNN). EDGE scales to large graphs (up to $\sim$4k nodes and $\sim$38k edges), however, the number of required diffusion steps and hence graph generation time increases linearly with the number of edges in the graph.
@@
To address the existing gaps in this field, we introduce ARROW-Diff (AutoRegressive RandOm Walk Diffusion), a novel iterative procedure designed for efficient, high-quality, and large-scale graph generation. To the best of our knowledge, ARROW-Diff is the first method to perform (discrete) diffusion on the level of random walks within a graph. Our method overcomes many limitations of existing graph generation approaches, including scalability issues, the independent treatment of edges, the need for post-processing, and the high computational complexity resulting from operations performed directly on the adjacency matrix. ARROW-Diff integrates two components: (1) A discrete diffusion model based on the order-agnostic autoregressive diffusion framework (OA-ARDM) (Hoogeboom et al., 2022) applied on random walks sampled from a real-world input graph, and (2) a GNN trained to predict the validity of the proposed edges in the generated random walks from component (1). The random walk-based diffusion in component (1) enables ARROW-Diff to capture the complex and sparse structure of graphs (Perozzi et al., 2014; Grover & Leskovec, 2016; Bojchevski et al., 2018), avoid any dense computation, and scale to large graphs, since the required number of diffusion steps is only equal to the random walk length. Moreover, due to the autoregressive nature of the implemented diffusion framework (OA-ARDM), ARROW-Diff overcomes the edge-independence limitation of methods like NetGAN (Bojchevski et al., 2018) and VGAE (Kipf & Welling, 2016b). On the other hand, the GNN in component (2) enables ARROW-Diff to select highly probable edges based on the topological structure learned from the original graph. It also helps to overcome the need for post-processing as in NetGAN (Bojchevski et al., 2018). ARROW-Diff builds the final graph by integrating the two components in an iterative process. This process is further guided by node degrees inspired by EDGE (Chen et al., 2023) to generate graphs with a similar degree distribution to the original graph. Using this iterative approach, we show that ARROW-Diff demonstrates superior or comparable performance to other baselines on multiple graph statistics. Furthermore, ARROW-Diff is able to efficiently generate large graphs (in this work, up to $\sim$20k nodes), such as the citation networks from McCallum et al. (2000); Sen et al. (2008); Pan et al. (2016), with significantly reduced (down to almost $50\%$) generation time compared to other baselines.
@@ 2 Related Work
These methods include graph generation approaches based on the Variational Autoencoder (VAE) (Kingma & Welling, 2013), like VGAE (Kipf & Welling, 2016b), GraphVAE (Simonovsky & Komodakis, 2018) and Graphite (Grover et al., 2019), which embed a graph $G$ into a continuous latent representation $\bm{z}$ using an encoder defined by a variational posterior $q_{\phi}(\bm{z}|G)$, and a generative decoder $p_{\theta}(G|\bm{z})$. These models are trained by minimizing the upper bound on the negative log-likelihood $\mathbb{E}_{q_{\phi}(\bm{z}|G)}[-\log p_{\theta}(G|\bm{z})]+\text{KL}[q_{\phi}(\bm{z}|G)||p(\bm{z})]$ (Kingma & Welling, 2013). However, due to their run time complexity of $\mathcal{O}(N^{2})$, VAE-based graph generation approaches are unable to scale to large graphs.
@@
The most scalable non-diffusion and sequential graph generation method so far is NetGAN (Bojchevski et al., 2018), which is based on the concept of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Specifically, it uses a Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) network to generate random walks. After training, the generated random walks are used to construct a score matrix from which the edges of the generated graph are sampled. The aforementioned approach generates edges in an independent manner, sacrificing the quality of the generated graphs and limiting their ability to reproduce some statistics of the original graphs such as triangle counts and clustering coefficient (Chanpuriya et al., 2021). In contrast, ARROW-Diff overcomes this problem by directly using the edges generated by an autoregressive diffusion process on the level of random walks. Other edge-dependent sequential approaches include GraphRNN (You et al., 2018) and GRAN (Liao et al., 2019). These methods iteratively generate the entries of a graph adjacency matrix one entry or one block of entries at a time. To overcome the long-term bottleneck issue of Recurrent Neural Networks (RNNs), Liao et al. (2019) propose to use a GNN architecture instead of an RNN, which utilizes the already generated graph structure to generate the next block, allowing it to model complex dependencies between each generation step. To satisfy the permutation invariance property of graphs, these methods require a node ordering scheme. Moreover, they are only able to scale to graphs of up to 5k nodes.
@@
To exploit the sparsity property of graphs, discrete diffusion-based graph generation models focus on diffusion in the discrete space i.e., on the level of the adjacency matrix (Haefeli et al., 2022; Vignac et al., 2023). Although these approaches generate high-quality graphs (Niu et al., 2020; Jo et al., 2022) and overcome the limitation of autoregressive models, they are restricted to small graph generation, like chemical molecules, because they perform prediction for every pair of nodes. Recently, Kong et al. (2023) developed GraphARM, an autoregressive graph generation approach based on diffusion in the node space. GraphARM demonstrates faster sampling time, since the number of diffusion steps required equals the number of nodes. Our method, however, requires only a number of diffusion steps equal to the random walk length, making it far more scalable than GraphARM. Aside from the method proposed in this work, the only diffusion-based approach able to scale to large graphs is EDGE (Chen et al., 2023). The forward diffusion process of EDGE is defined by successive edge removal until an empty graph is obtained. In the reverse diffusion process, only a fraction of edges are predicted between active nodes for which the degree changes during the forward diffusion process. This method generates graphs with a similar degree distribution to the original graph and has a decreased run time of $\mathcal{O}(T\max(M,K^{2}))$, where $M$ is the number of edges in a graph and $K$ is the number of active nodes. In this work, we introduce a random walk-based discrete diffusion approach to efficiently generate random walks as a first step for graph generation. The random walk-based diffusion implemented through the OA-ARDM (Hoogeboom et al., 2022) framework enables more efficient sampling than in other discrete diffusion models, with significantly reduced number of diffusion steps equal only to the random walk length. Additionally, a GNN component validates edges from the sampled random walks, enabling us to refine the generated graph and produce graphs of high quality.
@@ 3 Method
In this section, we introduce ARROW-Diff, an iterative procedure for efficient large-scale graph generation. We designed ARROW-Diff in a way to overcome the critical limitations of previous graph generation approaches and to enable it to scale to large graphs efficiently, without sacrificing the graph generation quality. Figure 1 provides an overview of ARROW-Diff, which leverages two models that are trained independently of each other on a training set of the edges from the original graph: (1) A random walk-based diffusion model based on the OA-ARDM (Hoogeboom et al., 2022) framework trained to generate random walks from the original graph. This involves first sampling of random walks and then masking nodes in the forward diffusion process, with the model predicting the original nodes in the reverse diffusion process. ARROW-Diff generates random walks from this learned diffusion model and uses the edges comprising these random walks as a first step of the graph generation process. (2) A GNN trained to capture the topology of the original graph. This component is used to predict the validity of the edges proposed by the first component. Our proposed method, ARROW-Diff, incorporates these two components in an iterative process guided by changes in the node degrees to generate the final graph. In the following, we provide background information on the OA-ARDM and its adaptation for random walk-based diffusion, followed by a detailed description of ARROW-Diff.
Figure 1: Overview of ARROW-Diff graph generation (inference) using a trained OA-ARDM (Hoogeboom et al., 2022) and a trained GNN. Iteratively, and starting from an empty graph, a diffusion model samples random walks from a set of start nodes. Then, a GNN classifies the proposed edges and filters out invalid ones. This procedure is repeated using a different, sampled set of start nodes guided by the change of node degrees w.r.t. the original graph.
@@
Our proposed graph generation method, ARROW-Diff, is outlined in Algorithm 2. ARROW-Diff generates new graphs similar to a single, given original graph $G=(V,E)$ through an iterative procedure of $L$ iterations: It starts with an empty graph $\hat{G}=(V,\hat{E}=\emptyset)$, which contains the same set of nodes $V$ as the original graph but no edges. To add edges to $\hat{G}$, ARROW-Diff initially samples one random walk for every node in $V$ and considers all edges in these random walks as edge proposals $\hat{E}_{\text{proposals}}$ for $\hat{G}$ (lines 6 and 7 in Algorithm 2). Next, and similar to the work of Liao et al. (2019), we sample valid edges from the proposed ones $\hat{E}_{\text{proposals}}$ using a Bernoulli distribution basen on the binary edge classification probabilities. These probabilities are calculated by taking the sigmoid of the dot-product of node embeddings predicted by the GNN component on $G=(V,\hat{E}\cup\hat{E}_{\text{proposals}},\bm{X})$ (line 8). This GNN model is trained on a perturbed version of a training set of edges from the original graph $G$ to perform edge classification. Specifically, the training set of edges is corrupted by deleting edges and inserting invalid (fake) edges. Inspired by the degree-guided graph generation process of Chen et al. (2023), we now sample nodes from $V$ using a Bernoulli distribution by considering each node $n\in V$ according to a success probability $p(n)=\frac{d_{n}}{\max(\bm{d})}$, where $\bm{d}:=\max(0,\bm{d}_{G}-\bm{d}_{\hat{G}})_{n}$ are the positive differences of node degrees $\bm{d}_{G}$ and $\bm{d}_{\hat{G}}$ from $G$ and $\hat{G}$. This set of sampled nodes is then used as start nodes for sampling the random walks in the next iteration. Hence, we modify the sampling procedure of Hoogeboom et al. (2022), which originally starts from a sequence with only masked tokens, by manually setting the first node of a random walk $\bm{x}$ to a specific node $n\in V$, i.e.
@@
Additionally, we use a restricted set of permutations $S_{D}^{(1)}:=\{\sigma\in S_{D}|\sigma(1)=1\}$, in which the order of the first element does not change after applying the permutation. To sample the remaining parts $\bm{x}_{2:D}$ of the random walk $\bm{x}$, we follow the sampling procedure of Hoogeboom et al. (2022) by starting at time step $t=2$ and using $\sigma\sim\mathcal{U}(S_{D}^{(1)})$. Our method can generate directed and undirected graphs. To generate undirected graphs, we suggest adding all reverse edges $\{(n_{j},n_{i})|(n_{i},n_{j})\in\hat{E}_{\text{proposals}}\}$ to the edge proposals $\hat{E}_{\text{proposals}}$ (line 7), and to sample edges from $\hat{E}\cup\hat{E}_{\text{proposals}}$ in a way to obtain undirected edges in $\hat{E}_{\text{valid}}$ (line 9).

### v2

# Random Walk Diffusion for Efficient Large-Scale Graph Generation

@@ 1 Introduction
Recently, deep learning-based approaches have gained attention for overcoming the limitations of traditional graph generation methods by learning the complex topology of real-world graphs. These approaches typically involve an encoder that learns a dense representation of the graph, a sampler that generates samples from this representation, and a decoder which restores the sampled representation into a graph structure (Zhu et al., 2022). Different decoding strategies have been proposed, including sequential decoders that generate the graph step-by-step (You et al., 2018; Liao et al., 2019), and one-shot generators that can create the entire graph in a single step (Kipf & Welling, 2016b; Simonovsky & Komodakis, 2018; Grover et al., 2019). While sequential decoders struggle to model long-term dependencies and require a node-ordering process, one-shot generators suffer from scalability issues and compromised quality due to the independent treatment of edges (Chanpuriya et al., 2021), making these methods suboptimal for large-scale graph generation (Guo & Zhao, 2022).
@@
An even more recent body of work in graph generation are diffusion-based probabilistic models, inspired by non-equilibrium thermodynamics and first introduced by Sohl-Dickstein et al. (2015). Briefly, diffusion models consist of two processes: A forward diffusion process which gradually corrupts input data until it reaches pure noise, and a reverse diffusion process which learns the generative mechanism of the original input data using a neural network. Diffusion-based methods for graph generation can be divided into two main categories. The first one includes methods that implement diffusion in the continuous space e.g., by adding Gaussian noise to the node features and graph adjacency matrix (Niu et al., 2020; Jo et al., 2022). This form of diffusion however makes it difficult to capture the underlying structure of graphs since it destroys their sparsity pattern (Vignac et al., 2023). The second category includes methods that are based on diffusion in the discrete space (Haefeli et al., 2022; Vignac et al., 2023; Chen et al., 2023) by performing successive graph modifications e.g., adding or deleting edges/nodes or edge/node features. Diffusion-based graph generation methods do not suffer from long-term memory dependency which makes them advantageous over autoregressive (sequential) methods. However, many approaches found in the literature are only designed for small graphs (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2023). One of the recent diffusion-based graph generation approaches is EDGE (Chen et al., 2023), which uses a combination of a discrete diffusion model and a Graph Neural Network (GNN). EDGE scales to large graphs (up to $\sim$4k nodes and $\sim$38k edges), however, the number of required diffusion steps and hence graph generation time increases linearly with the number of edges in the graph.
@@
To address the existing gaps in this field, we introduce ARROW-Diff (AutoRegressive RandOm Walk Diffusion), a novel iterative procedure designed for efficient, high-quality, and large-scale graph generation. To the best of our knowledge, ARROW-Diff is the first method to perform (discrete) diffusion on the level of random walks within a graph. Our method overcomes many limitations of existing graph generation approaches, including scalability issues, the independent treatment of edges, the need for post-processing, and the high computational complexity resulting from operations performed directly on the adjacency matrix. ARROW-Diff integrates two components: (1) A discrete diffusion model based on the order-agnostic autoregressive diffusion framework (OA-ARDM) (Hoogeboom et al., 2022) applied on random walks sampled from a real-world input graph, and (2) a GNN trained to predict the validity of the proposed edges in the generated random walks from component (1). The random walk-based diffusion in component (1) enables ARROW-Diff to capture the complex and sparse structure of graphs (Perozzi et al., 2014; Grover & Leskovec, 2016; Bojchevski et al., 2018), avoid any dense computation, and scale to large graphs, since the required number of diffusion steps is only equal to the random walk length. Moreover, due to the autoregressive nature of the implemented diffusion framework (OA-ARDM), ARROW-Diff overcomes the edge-independence limitation of methods like VGAE (Kipf & Welling, 2016b) and NetGAN (Bojchevski et al., 2018). On the other hand, the GNN in component (2) enables ARROW-Diff to select highly probable edges based on the topological structure learned from the original graph. It also helps to overcome the need for post-processing as in NetGAN (Bojchevski et al., 2018). ARROW-Diff builds the final graph by integrating the two components in an iterative process. This process is further guided by node degrees inspired by EDGE (Chen et al., 2023) to generate graphs with a similar degree distribution to the original graph. Using this iterative approach, we show that ARROW-Diff demonstrates superior or comparable performance to other baselines on multiple graph statistics. Furthermore, ARROW-Diff is able to efficiently generate large graphs (in this work, up to $\sim$20k nodes and $\sim$60k edges), such as the citation networks from McCallum et al. (2000); Sen et al. (2008); Pan et al. (2016), with significantly reduced (down to almost $50\%$) generation time compared to other baselines.
@@ 2 Related Work
These methods include graph generation approaches based on the Variational Autoencoder (VAE) (Kingma & Welling, 2013) model, like VGAE (Kipf & Welling, 2016b), GraphVAE (Simonovsky & Komodakis, 2018) and Graphite (Grover et al., 2019), which embed a graph $G$ into a continuous latent representation $\bm{z}$ using an encoder defined by a variational posterior $q_{\phi}(\bm{z}|G)$, and a generative decoder $p_{\theta}(G|\bm{z})$. These models are trained by minimizing the upper bound on the negative log-likelihood $\mathbb{E}_{q_{\phi}(\bm{z}|G)}[-\log p_{\theta}(G|\bm{z})]+\text{KL}[q_{\phi}(\bm{z}|G)\,||\,p(\bm{z})]$ (Kingma & Welling, 2013). However, due to their run time complexity of $\mathcal{O}(N^{2})$, VAE-based graph generation approaches are unable to scale to large graphs.
@@
One of the most scalable non-diffusion and sequential graph generation methods is NetGAN (Bojchevski et al., 2018), which is based on the concept of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Specifically, it uses a Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) network to generate random walks. After training, the generated random walks are used to construct a score matrix from which the edges of the generated graph are sampled. The aforementioned approach generates edges in an independent manner, sacrificing the quality of the generated graphs and limiting their ability to reproduce some statistics of the original graphs such as triangle counts and clustering coefficient (Chanpuriya et al., 2021). In contrast, ARROW-Diff overcomes this problem by directly using the edges generated by an autoregressive diffusion process on the level of random walks. Other edge-dependent sequential approaches include GraphRNN (You et al., 2018), GRAN (Liao et al., 2019), and BiGG (Dai et al., 2020). These methods iteratively generate the entries of a graph adjacency matrix one entry or one block of entries at a time. To overcome the long-term bottleneck issue of Recurrent Neural Networks (RNNs), Liao et al. (2019) propose to use a GNN architecture instead of an RNN, which utilizes the already generated graph structure to generate the next block, allowing it to model complex dependencies between each generation step. To satisfy the permutation invariance property of graphs, these methods require a node ordering scheme. Moreover, GRAN and GraphRNN are only able to scale to graphs of up to 5k nodes.
@@
To exploit the sparsity property of graphs, discrete diffusion-based graph generation models focus on diffusion in the discrete space, i.e., on the level of the adjacency matrix (Haefeli et al., 2022; Vignac et al., 2023). Although these approaches generate high-quality graphs (Niu et al., 2020; Jo et al., 2022) and overcome the limitation of autoregressive models, they are restricted to small graph generation, like chemical molecules, because they perform predictions for every pair of nodes. Recently, Kong et al. (2023) developed GraphARM, an autoregressive graph generation approach based on diffusion in the node space. GraphARM demonstrates faster sampling time, since the number of diffusion steps required equals the number of nodes. Our method, however, requires only a number of diffusion steps equal to the random walk length, making it far more scalable than GraphARM. Aside from the method proposed in this work, the only diffusion-based approach able to scale to large graphs is EDGE (Chen et al., 2023). The forward diffusion process of EDGE is defined by successive edge removal until an empty graph is obtained. In the reverse diffusion process, only a fraction of edges are predicted between active nodes for which the degree changes during the forward diffusion process. This method generates graphs with a similar degree distribution to the original graph and has a decreased run time of $\mathcal{O}(T\max(M,K^{2}))$, where $M$ is the number of edges in a graph and $K$ is the number of active nodes. In this work, we introduce a random walk-based discrete diffusion approach to efficiently generate random walks as a first step for graph generation. The random walk-based diffusion process implemented through the OA-ARDM (Hoogeboom et al., 2022) framework enables more efficient sampling than in other discrete diffusion models, with significantly reduced number of diffusion steps equal only to the random walk length. Additionally, a GNN component validates edges from the sampled random walks, enabling us to refine the generated graph and produce graphs of high quality.
@@ 3 Method
In this section, we introduce ARROW-Diff, an iterative procedure for efficient large-scale graph generation. We designed ARROW-Diff in a way to overcome the critical limitations of previous graph generation approaches and to enable it to scale to large graphs efficiently, without sacrificing the graph generation quality. Figure 1 provides an overview of ARROW-Diff, which leverages two models that are trained independently of each other on a training set of edges from the original graph: (1) A random walk-based diffusion model based on the OA-ARDM (Hoogeboom et al., 2022) framework trained to generate random walks from the original graph. This involves first sampling of random walks and then masking nodes in the forward diffusion process, with the model predicting the original nodes in the reverse diffusion process. (2) A GNN trained to capture the topology of the original graph. Our proposed method, ARROW-Diff, incorporates these two components in an iterative process in which (1) the diffusion model generates random walks, proposing potential edges, and (2) the GNN evaluates and filters the edges proposed by the OA-ARDM based on learned topological features. This process is further guided by changes in the node degrees to generate the final graph. In the following, we provide background information on the OA-ARDM and its adaptation for random walk-based diffusion, followed by a detailed description of ARROW-Diff.
Figure 1: Overview of ARROW-Diff graph generation (inference) using a trained OA-ARDM (Hoogeboom et al., 2022) and a trained GNN. Iteratively, and starting from an empty graph, a diffusion model samples random walks from a set of start nodes. Then, a GNN classifies the proposed edges and filters out invalid ones. This procedure is repeated $L$ times using a different set of sampled start nodes guided by the change of node degrees with respect to the original graph.
@@
Our proposed graph generation method, ARROW-Diff, is outlined in Algorithm 2. ARROW-Diff generates new graphs similar to a single, given original graph $G=(V,E)$ through an iterative procedure of $L$ iterations: It starts with an empty graph $\hat{G}=(V,\hat{E}=\emptyset)$, which contains the same set of nodes $V$ as the original graph but no edges. To add edges to $\hat{G}$, ARROW-Diff initially samples one random walk for every node in $V$ and considers all edges in these random walks as edge proposals $\hat{E}_{\text{proposals}}$ for $\hat{G}$ (lines 6 and 7 in Algorithm 2). Next, and similar to the work of Liao et al. (2019), we sample valid edges from the proposed ones $\hat{E}_{\text{proposals}}$ using a Bernoulli distribution based on the binary edge classification probabilities. These probabilities are calculated by taking the sigmoid of the dot-product of node embeddings (for node-pairs inside $\hat{E}\cup\hat{E}_{\text{proposals}}$) predicted by the GNN component on $G=(V,\hat{E}\cup\hat{E}_{\text{proposals}},\bm{X})$ (line 8). This GNN model is trained on a perturbed version of a training set of edges from the original graph $G$ to perform edge classification. Specifically, the training set of edges is corrupted by deleting edges and inserting invalid (fake) edges. Inspired by the degree-guided graph generation process of Chen et al. (2023), we now sample nodes from $V$ using a Bernoulli distribution by considering each node $n\in V$ according to a success probability $p(n)=\frac{d_{n}}{\max(\bm{d})}$, where $\bm{d}:=\max(0,\bm{d}_{G}-\bm{d}_{\hat{G}})_{n}$ are the positive differences of node degrees $\bm{d}_{G}$ and $\bm{d}_{\hat{G}}$ from $G$ and $\hat{G}$. This set of sampled nodes is then used as start nodes for sampling the random walks in the next iteration. Hence, we modify the sampling procedure of Hoogeboom et al. (2022), which originally starts from a sequence with only masked tokens, by manually setting the first node of a random walk $\bm{x}$ to a specific node $n\in V$, i.e.
@@
Additionally, we use a restricted set of permutations $S_{D}^{(1)}:=\{\sigma\in S_{D}|\sigma(1)=1\}$, in which the order of the first element does not change after applying the permutation. To sample the remaining parts $\bm{x}_{2:D}$ of the random walk $\bm{x}$, we follow the sampling procedure of Hoogeboom et al. (2022) by starting at time step $t=2$ and using $\sigma\sim\mathcal{U}(S_{D}^{(1)})$. ARROW-Diff can generate directed and undirected graphs. To generate undirected graphs, we suggest adding all reverse edges $\{(n_{j},n_{i})|(n_{i},n_{j})\in\hat{E}_{\text{proposals}}\}$ to the edge proposals $\hat{E}_{\text{proposals}}$ (line 7), and to sample edges from $\hat{E}\cup\hat{E}_{\text{proposals}}$ in a way to obtain undirected edges in $\hat{E}_{\text{valid}}$ (line 9).